Please help me solve this problem?
Twenty-five boys and fifteen girls joined the Mathematics Club. What percent of the members were girls?

Answer by lenny460(1073)

Please help me solve this problem?
Twenty-five boys and fifteen girls joined the Mathematics Club. What percent of the members were girls?
25 boys and 15 girls.
Therefore:
25 + 15 = 100%
25 + 15 = 40
40 = 100%
10% = 40/10 = 4 students
5% = 2 students
2 1/2% = 1 student

25 boys = 4 * 6 + 1
10 * 6 = 60 + 2 1/2
Therefore the percentage of boys is:
60 + 2 1/2 = 62 1/2%

15 girls = 4 * 3 + 2 + 1
4 * 3 + 2 + 1 = 12 + 2 + 1 = 15
10 * 3 + 5 + 2 1/2
30 + 5 + 2 1/2 = 37 1/2%
The percentage of girls is:
37 1/2%



Lennox Obuong
Algebra Tutor